If u be a borel measure on a metric space X and suppose there exists a sequence Gn of open sets, X=UnGn and for all n , u(Gn) < Infinity then u is regular.


1
Expert's answer
2020-10-15T18:35:47-0400



Let M= collection of all borel measurable subsets "Y" of X such that

"\mu(Y)=sup\{\mu(K)|K\subset X,K \ closed\}=inf\{\mu(U)|U\subset X,U \ open\}". We will show M is "\sigma-" algebra containing all closed sets and hence all open sets. This will show "\mu" is regular.

Now let "Y\in M." Then, for "\epsilon >0," there is "U" open and "K" closed such that, "\mu(U)<\mu(Y)+\epsilon""\mu(K)>\mu(Y)-\epsilon". Hence "\mu(Y^c)> \mu(U^c)+\epsilon" and "\mu(Y^c)< \mu(K^c)-\epsilon" Hence M is closed under complementation. Now let "Y_i" be a countable collection. Then there exists "U_i" open and "K_{i}" closed such that "\mu(K_i)>\mu(Y_i)-\epsilon /2^i" and "\mu(U_i)<\mu(Y_i)+\epsilon /2^i" Let "K=\cap K_i" and "U=\cup U_i" . Then since its given the measure is "\sigma" finite, "\mu(U)=lim_{n\rightarrow \infty}\mu(U_i)< \mu(Y)+\Sigma_{i=1}^{\infty} \ \epsilon/2^i" "=\mu(Y)+\epsilon". Similarly "\mu(K)>\mu(Y)-\epsilon." Hence M is closed under countable union. Now we have to show that it contains all closed sets. Let "Y" be closed. Let "U_n=\{x\in X| d(y,x)<1/n \ \textrm{for some} \ y \in Y \}." Then "U_n" open and "U_1\supset U_2\supset U_3\cdots" Also as "Y" is closed, so for any "x\notin Y" "d(x,Y)=m>0. \Rightarrow m>1/n" for some "n." Hence "\cap_i U_i =Y" Hence "\mu(Y)= lim \ \mu(U_n)=" inf "\mu(U_n)" "\geq \textrm{inf} \ \mu(U) \ \forall \ U\supset Y." . But RHS "\geq \mu(Y)" and hence the equality. For the other side we note "Y"is a closed set and hence is the supremum. Hence all closed sets are in M. Hence done.